A particle starts from origin at t=0 witha vel of 8j m/s and moves in x-y plane with const.acclrtn of (4i+2j)m/s2.At the instant the particle's x-coordinate is 29m,what are


1)its y-coordinate?


2)Its speed?




 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

for motion in x direction

Ux=0

a= 4 m/sec2

Sx= 29 m

Sx=Ux t+1/2 a t2

29 =0+1/2 * 4 * t2

t2 =29/2  ........................1

Vx= Ux +at

  Vx= 4t ......................2

Motion in y direction

Sy = Uyt+1/2 a t2

Sy =8 t +1/2 * 2* t2

    =8t +t2               Answer

put value of t from equation 1 and find Sy

Vy =Uy + at

     =8  +2t   put value of t from equation 1 and find Vy       

 

now speed at this point =( Vx2 +Vy2)1/2                 Answer


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin